OpenStudy (anonymous):

(3/2p-2)-1=(1/p-1)+2 How do I solve for p??

OpenStudy (anonymous):

p=4+ 22‾‾‾√ 3,4− 22‾‾‾√ 3 p≈2.89680525,−0.23013859

OpenStudy (anonymous):

Let's solve your equation step-by-step. 32p−2−1=1p−1+2 32p−3=p+1p Step 1: Multiply both sides by p. 32p2−3p=p+1 32p2−3p−(p+1)=p+1−(p+1)(Subtract p+1 from both sides) 32p2−4p−1=0 p=−b±b2−4ac2a(Use quadratic formula with a=1.5, b=-4, c=-1) p=−(−4)±(−4)2−4(1.5)(−1)2(1.5) p=4±223 p=2.896805253274477,−0.23013858660780992 Check answers. (Plug them in to make sure they work.) p=2.896805(Works in original equation) p=−0.230139(Works in original equation) Answer: p=2.896805 or p=−0.230139